(ii)       Eliminate R from equations (i) and (ii) in 4 above and obtain an equation connecting E, V, I and r. Identify the independent and dependent variable in this equation and write it in the y = mx + b form.                 

eq 1E=I(R+r)

eq 2 V=IR
Answer #1

E = I (Rt) - W V=IR - (ii) Expanding eq ci) E=IRtI8 - (ii) from cing and iii) E = V + Ir -) [E = x3 + v Here independent vari